# Enabling Quick, Accurate Crowdsourced Annotation for Elevation-Aware Flood Extent Mapping

### v1

@@ 4 Evaluation
In this evaluation, we study how FloodTrace is able to enable crowdsourced annotation through an experimental user study. The setup of this study is explained in subsection 4.1. We use the results of this study to assess the efficiency of our semi-automatic tools, how the insight given by these and 3D rendering affect annotation accuracy, and the value of our aggregate visualization and correction strategy. To do this, we first analyze the accuracy and annotation speed of participants (subsection 4.2). We then use participant annotations as training data for flood detection models and assess model quality (subsection 4.3).
@@ 4.1 User Study Setup
The dataset used for annotation for the study consists of flooded regions of North Carolina during Hurricane Matthew in 2016, with high-resolution aerial imagery from the National Oceanic and Atmospheric Administration National Geodetic Survey [47] and corresponding DEM data from the University of North Carolina Libraries [12]. All data was resampled to a resolution of 2 square meters per pixel, which is common in this domain [37, 53, 68]. Eight separate regions were chosen with dimensions between 4104 $\times$ 1856 pixels and 6472 $\times$ 3136 pixels, and these regions were split into quadrants for smaller workloads for participants of the study.
297 graduate students of a machine learning course volunteered to participate in this study, of which 266 followed through with final submissions. Before beginning annotation work, participants watched an instructional video explaining how to use all of the features of FloodTrace and how to determine whether to mark areas as dry or flooded. They were then given access to a web application which they could visit at any time within the following two weeks to complete the work assigned to them. On the application, all interactions made by the students were logged, and these logs were submitted along with output annotations on completion. Each student was asked to complete five annotation tasks, with each task requiring annotating at least $60\%$ of a quadrant of one of our chosen regions as flooded or dry, leaving any remaining pixels as unlabeled. These tasks were created for two specific experiments, in addition to providing a reasonable number of annotations and interaction logs to draw insight from.
@@
The second experiment was meant to test the effectiveness of the 3D mesh in improving participants’ insight and accuracy. This required users to perform two annotation tasks on the same quadrant, although a different one than used in the first three tasks. First, they would annotate the quadrant while restricted to only a 2D view of the aerial imagery, then they would annotate the same quadrant with the 3D mesh and full 3D view interactivity. Both tasks for this experiment had full access to brush and semi-automatic tools.
@@ 4.2 User Study Analysis
Because the goals of R1 and R2 are to increase annotation speed and accuracy, these are the metrics we study. We measure annotation speed simply as the time taken to complete an annotation. We measure annotation accuracy by comparing participant annotations against reference annotations that were manually labeled either by our domain expert collaborators, or members of our team after being trained. To avoid penalizing unlabeled pixels, we only consider pixels that were labeled as flooded or dry in both the participant and the reference annotation when measuring accuracy. The formula we use to compute accuracy is shown below where TF and TD are true flooded and true dry (pixels with the same label in participant and reference), and FF and FD are false flooded and false dry (pixels with opposite labels in participant and reference):
@@
An accuracy percentage of 0% here means all participant-labeled pixels disagreed with all corresponding labels in the reference, 100% means all user-labeled pixels agreed with corresponding labels in the reference, and 50% means half of the user-labeled pixels agreed with corresponding labels in the reference. Note a score of 50% can be achieved by labeling pixels completely randomly since there is a 50% chance of guessing a flooded or dry label correctly.
Impact of Annotation Tools
In order to study how the usage of different tools affected these metrics, we grouped annotations by their tool usage using the activity logs submitted with participants’ annotations. Before analysis, we first conducted data cleaning to remove potential spoilers from the dataset, as we found some participants to misunderstand the labeling task or leave the application open without progress for extremely long periods. We computed the means and standard deviations for annotation completion time and accuracy, and used these to compute a threshold at which to discard annotations. We chose a standard z-score of 3 as our cutoff point (meaning 3 standard deviations from the mean, or falling outside of the 99.7% confidence interval), and were left with thresholds of 17.6 hours for annotation time and 55.4% for accuracy. Because the variance of annotation times was much higher than the variance of accuracies, the time threshold could likely be more aggressive; the median completion time was only 32 minutes, so it seems probable that times above a few hours would be the result of idling on the page. We investigated using z-scores of 1 and 2 to create stricter time thresholds (corresponding to 7.0 and 12.3 hours respectively), but found that these led to the same statistically significant conclusions as using a z-score of 3, only more pronounced, so we present here the results using the more inclusive 17.6 hour cutoff. With our time and accuracy thresholds, we discard a total of 40 annotations and leave 1,281 activity logs for analysis. Using these, we create four groups based on tool usage as described below:
@@
  • Elevation Guided: Annotations where elevation-guided tools (segmentation and both BFS selections) combined were used to label $>50\%$ of all labeled pixels
In order, these groups ended up with sizes of 787, 134, 346, and 494 annotations. We find that, even though the instructional video emphasized that participants should rely mostly on the elevation-guided tools, users gravitated more towards the more simple and familiar brush tool. This likely shows an example of algorithm aversion [14, 13], where users are proven to prefer human decision-making rather than algorithmic tools even when presented with algorithms that outperform humans. We acknowledge that this could be fixed in future deployments by forcing some usage of the elevation-guided tools before the brush tool is unlocked, but we leave more in-depth approaches to overcoming algorithm aversion as future work.
Figure 8: Box and whisker plot showing the accuracy of the brush, BFS, segmentation, and elevation guided groups.
Brush BFS Segmentation Elevation Guided
@@
Table 1: Mean accuracies and p-values from t-tests against the brush group. We find all three groups that rely on elevation-guided tools to be more accurate than the group using the manual tool. All comparisons are statistically significant except the BFS group, likely because the group had a much smaller sample size.
@@
We present annotation accuracy results for these groups in a box and whisker plot in Figure 8. We provide the mean accuracies in Table 1, along with Welch’s t-test results to test for statistical significance in difference against the brush group. From these results, we find that those who relied on elevation-guided tools were able to produce more accurate annotations than those who relied on the brush tool, which is guided solely by user insight. This shows that both elevation-guided tools are effective for improving accuracy (and thus training data quality), even for untrained crowdworkers. We additionally ran t-tests for the differences between the BFS, segmentation, and elevation-guided groups, but found no statistically significant outcomes.
@@
Figure 9: Box and whisker plot showing the time taken for annotations in the brush, BFS, segmentation, and elevation guided groups. We find use of the segmentation tool to drastically reduce the time taken to complete an annotation, with brush and BFS groups having median times of 39.97 and 38.18 minutes while segmentation and elevation-guided groups took only 17.45 and 21.65 minutes.
@@
We present annotation speed results in a box and whisker plot in Figure 9. This plot shows that annotations relying on the segmentation tool were much faster than those that did not, with the median time taken of the segmentation group being 2.29×2.29$\times$2.29 × and 2.19×2.19$\times$2.19 × smaller than the medians of the brush and BFS groups respectively. We ran t-tests again to measure the significance of the differences in time taken between groups, and found the differences between segmentation and brush, segmentation and bfs, and elevation-guided and brush to be significant with p-values < 0.00, with other group differences not significant except elevation-guided and BFS at 0.036. These results clearly support our segmentation tool as a method of quick annotation while preserving high accuracy, and show it outperforms the BFS tool in efficiency. Importantly, these results show that even untrained crowdworkers can use this tool effectively.
Tool Time Taken Accuracy
r P-Value r P-Value
BFS 0.022 0.218 0.053 0.029
Segmentation -0.137 0.000 0.034 0.114
Brush 0.111 $0$0.065 0.010
Table 2: Pearson correlation coefficients r and p-values for relationships between tool usage and our accuracy and speed metrics. P-values under 0.05 indicate statistical significance. Coefficients close to 0 indicate no correlation, while values closer to 1 and -1 indicate positive and negative correlations. All correlations are significant except time taken with BFS tool and accuracy with segmentation tool.
To further explore how the annotation tools affect accuracy and speed, we additionally compute the Pearson correlation coefficient r between each tool’s usage and our metrics. Tool usage is computed as the ratio of the number of pixels annotated by a tool divided by the total pixels labeled in that annotation. This data is presented in Table 2. The correlation coefficient r is measured in $[-1,1]$ where -1 is the strongest negative correlation, 1 is the strongest positive correlation, and 0 is no correlation.
Results show that the segmentation tool has a strong negative correlation with time taken, while brush had a strong positive correlation and BFS had no significant correlation. This shows that the more a participant used the segmentation tool, the faster they were able to complete their annotation. With regards to accuracy, usage of the BFS tool correlated with higher accuracy, meaning increased usage led to increased accuracy in the resulting annotation, and opposite for the brush tool. The correlation with accuracy was also positive for the segmentation tool, but it was smaller and not statistically significant. This shows that, while annotations that utilized the segmentation tool were accurate (i.e. those in the segmentation group), increasing usage of the tool did not always mean increased accuracy. Although the segmentation tool is the quickest way to annotate accurately, annotations are most accurate when it is supplemented with the BFS or brush tools to fill in gaps and make small corrections.
Impact of 3D View
We also assess how providing users with a 3D rendering affects their accuracy, using the two annotation tasks from the second experiment in subsection 4.1. As a note, participants’ accuracy in these tasks was higher than the overall average, stemming from the fact that participants completed these annotation tasks last, and their accuracy with the tools improved through learning over time. For this experiment we form two groups, flat, consisting of the annotation tasks restricted to the flat view, and 3D, consisting of the subsequent annotation tasks with access to the 3D rendering. Using the same accuracy and time thresholds as above, we get groups of 262 and 250 annotations for flat and 3D respectively. The mean accuracies for flat and 3D were 91.90 and 92.94, and the p-value for the t-test between them was 0.094, which is not significant at a 95% confidence interval. We hypothesize that the difference in accuracy was narrowed by the use of elevation-guided tools, which were able to compensate for the insight lost from not being able to view the 3D rendering. We tested this by rerunning the analysis while restricting the flat and 3D groups to only include annotations in the brush group. These new flat and 3D groups included 150 and 144 annotations respectively, had means of 90.96 and 92.69, and the p-value computed between the difference in their accuracies was 0.058. While these results are still not significant, the difference in accuracy did become larger, and the p-value lower. In the future, we hope that extending this experiment with a larger sample size could lead to more conclusive results.
@@ 4.3 Crowdsourced Annotations for ML Model Training
In this section, we show how FloodTrace can enable researchers to use crowdsourced annotations as high-quality training data through an aggregation and correction process, which we detail in subsubsection 4.3.1. We propose this step to replace the extremely time-consuming annotation step of researchers’ workflows. To assess the training data quality of the results from aggregation and correction, we use study participants’ annotations as training data for deep learning models and compare them against models trained with reference annotations subsubsection 4.3.2.
@@ 4.3.1 Example Aggregation and Correction
As detailed in subsection 3.3, our system offers the capability to visualize aggregated crowdsourced annotations and fix inaccuracies. We conducted an example for this on two randomly selected regions, collecting all the annotations for those regions from our user study. We purposefully did not use any accuracy, time, or tool usage cutoffs when creating these sets of annotations in order to ensure there would be inaccuracies in the aggregate data. For each region, there was an average of 45 annotations per pixel. We input these sets to the application frontend, and one of our trained users utilized our aggregate and uncertainty views to correct each region. On average, $4\%$ of each region’s pixels were corrected. This process was timed, and it took an average of 24 minutes to make each region’s correction annotation. We compare this to the average time it took our trained users to create reference annotations in our application, which was 116 minutes per region. As training robust flood detection models requires many different annotated regions, this demonstrates how the crowdsourcing and correction workflow can save researchers huge amounts of time compared to the typical annotation process.
@@ 4.3.2 Prediction Setup and Results
In order to use our crowdsourced annotations (corrected and uncorrected) to train our models, we preprocess them by turning each group of annotations into a training set of soft labels. This is done by computing the flood and dry scores for each pixel, where the flood score is the sum of the number of times it was annotated as flooded divided by the number of times it was annotated as either flooded or dry, with dry scores using the same formula. We ignore unlabeled pixels in each annotation to compute these scores.
@@
Table 3: Performance metrics for models trained using uncorrected, corrected, and reference labels as ground truths.
We initialized three models, using the architecture described, and trained each of them in the same two regions. Each model used a different set of labels as its respective ground truths during training, with one model using the uncorrected crowdsourced labels, one using the corrected crowdsourced labels, and one using our reference labels. We then use these three trained models to predict flooded and dry labels for an unseen test region and compute quality metrics by comparing them with the reference annotation for that region. We show these results for each group in Table 3. We find that, with the corrected crowdsourced labels, we are able to achieve around equal results (measured in accuracy and f-score) to our expert-labeled references and improve dramatically over the uncorrected labels.
@@ 5 Conclusion and Limitations
We have proposed FloodTrace, a novel web-based framework for quick and accurate annotation of flooded regions that better enables crowdsourcing. This framework was built directly from requirements gathered from researchers in flood extent mapping, and incorporated their feedback continuously through development. Our application utilizes a region’s DEM elevation data to provide users insight in 3D and guide users with semi-automatic annotation tools, improving their accuracy. This application has already been utilized for the creation of training data by domain experts for cutting-edge flood extent mapping models [54]. The application includes a novel annotation method that uses topological data analysis to greatly outperform the state-of-the-art elevation-guided tool in efficiency. Our experimental user study shows that the benefits of our tools can be found even for untrained users. With our method for aggregating and correcting groups of annotations, researchers are able to use crowdsourced annotations as training data with equal performance to fully expert-labeled data. This improves their workflows greatly, replacing an extremely time-consuming annotation process with a much quicker aggregation and correction step.
@@
Our work is not without limitations, though. For example, our work only supports labeling for two classes of pixels, flooded or dry. Selecting features of the data with more specific labels, such as permanent water bodies, is useful for some flood detection models. Future work could tackle extending elevation-guided tools to annotation tasks with more informative labels than simply flooded or dry. There are also limitations with regards to the use of DEM data. As with the machine learning algorithms that rely on DEM data, our system requires that the DEM data and flooded aerial imagery are collected at similar dates. If not, deviations between elevation at the time of flooding and time of DEM collection could lead to annotation error. In addition, it is required that the DEM data precisely matches the region of aerial imagery being annotated, which necessitates pre-processing by the researcher. For the study discussed, we downsampled DEM data to match the spatial resolution of the aerial imagery, and have not investigated the effect of using DEM data and aerial imagery with differing resolutions, which could lead to conflict in annotation. Last, the use of DEM data means that objects above the ground level are absent. This is important for labeling areas with tree canopies and buildings, as we wish to consider them flooded if the ground level is flooded. However, for some regions, such as heavily urban environments with many buildings, the presence of these objects could affect water propagation enough to cause the semi-automatic tools to label things incorrectly, for example causing the BFS to falsely label downstream pixels as flooded when the water actually stopped at a building. In the future, it would be interesting to consider semi-automatic tools that utilize digital surface model (DSM) data, taking into account above-ground structures, for use cases on more urban regions.

### v2

@@ 4 Evaluation
In this evaluation, we study FloodTrace through an experimental user study explained in subsection 4.1. We use the results of this study to assess the efficiency of our topological segmentation tool and the value of our aggregate visualization and correction strategy. To do this, we first analyze the accuracy and annotation speed of participants (subsection 4.2). We then use participant annotations as training data for flood detection models and assess model quality (subsection 4.3).
@@ 4.1 User Study Setup
The dataset used for the study consists of flooded regions of North Carolina during Hurricane Matthew in 2016, with high-resolution aerial imagery from the National Oceanic and Atmospheric Administration National Geodetic Survey [47] and corresponding digital elevation model (DEM) data from the University of North Carolina Libraries [14]. All data was resampled to a resolution of 2 square meters per pixel, which is common in this domain [38, 53, 67]. Eight separate regions were chosen with dimensions between 4104 $\times$ 1856 pixels and 6472 $\times$ 3136 pixels, and these regions were split into quadrants for smaller workloads for participants of the study.
297 graduate students of a machine learning course volunteered to participate in this study, of which 266 followed through with final submissions. Before beginning annotation work, participants watched an instructional video explaining how to use all of the features of FloodTrace and how to determine whether to mark areas as dry or flooded. They were then given access to a web application which they could visit at any time within the following two weeks to complete the work assigned to them. On the application, all interactions made by the students were logged, and these logs were submitted with output annotations on completion. Each student was asked to complete five annotation tasks, with each task requiring annotating at least $60\%$ of a quadrant of one of our chosen regions as flooded or dry, leaving any remaining pixels as unlabeled. These tasks were created for two specific experiments, in addition to providing a reasonable number of annotations and interaction logs to draw insight from.
@@
The second experiment was meant to test the effectiveness of the 3D mesh in improving participants’ insight and accuracy. This required users to perform two annotation tasks on the same quadrant, although different form the one used in the first experiment. In this experiment, students would annotate while restricted to only a 2D view of the aerial imagery, then they would annotate the same quadrant with the 3D mesh and full 3D view interactivity. Both tasks for this experiment had full access to brush and semi-automatic tools.
@@ 4.2 User Study Analysis
In this section, we analyze the results of our user study to draw conclusions about how the usage of different annotation tools (brush, BFS, and segmentation) affected annotation speed and accuracy. Specifically, we test the hypotheses that elevation-guided tools increase accuracy, while the segmentation tool decreases the time needed for annotation. We also assessed how providing users with a 3D rendering affected their accuracy, using the two annotation tasks from the second experiment in subsection 4.1, but found no statistically significant results so omit these results for space. In subsubsection 4.2.1 we first define our metrics and describe how the data was preprocessed. Next, to compare the tools against each other and quantify the differences in speed and accuracy, we divide annotations into groups depending on which tool was used for the majority of labeling and compute differences between groups in subsubsection 4.2.2 . For all statistical tests, we select a 95% confidence interval for significance, then use the Benjamini-Hochberg (BH) procedure to correct for multiple testing by limiting the false discovery rate to 0.05 in each family of tests. Necessary p-value thresholds for significance computed by BH ($\alpha$) are presented with each table.
4.2.1 Tool Analysis Setup
Because the goals of R1 and R3 are to increase annotation speed and accuracy, these are the metrics we study. We measure annotation speed simply as the time taken to complete an annotation. We measure annotation accuracy by comparing participant annotations against reference annotations that were manually labeled either by our domain expert collaborators, or members of our team after being trained. To avoid penalizing unlabeled pixels, we only consider pixels that were labeled as flooded or dry in both the participant and the reference annotation when measuring accuracy. The formula we use to compute accuracy is shown below. TF and TD denote true flooded and true dry (pixels with the same label in participant and reference), and FF and FD are false flooded and false dry (pixels with opposite labels in participant and reference):
@@ 4.2.1 Tool Analysis Setup
An accuracy percentage of 0% here means all participant-labeled pixels disagreed with all corresponding labels in the reference, 100% means all user-labeled pixels agreed with corresponding labels in the reference, and 50% means half of the user-labeled pixels agreed with corresponding labels in the reference. Note that a score of 50% can be achieved by labeling pixels randomly since there is a 50% chance of guessing a flooded or dry label correctly.
Before statistical analysis, we first conducted data cleaning to remove potential spoilers from the dataset, as we found some participants to misunderstand the labeling task or leave the application open without progress for extremely long periods. We computed the means and standard deviations for annotation completion time and accuracy, and used these to compute a threshold at which to discard annotations. We chose a standard z-score of 3 as our cutoff point (meaning 3 standard deviations from the mean, or falling outside of the 99.7% confidence interval), and were left with thresholds of 17.6 hours for annotation time and 55.4% for accuracy. Because the variance of annotation times was much higher than the variance of accuracies, the time threshold could likely be more aggressive; the median completion time was only 32 minutes, so it seems probable that times above a few hours would be the result of idling on the page. We investigated using z-scores of 1 and 2 to create stricter time thresholds (corresponding to 7.0 and 12.3 hours respectively), but found that these led to the same statistically significant conclusions as using a z-score of 3, only more pronounced, so we present here the results using the more inclusive 17.6 hour cutoff. With our time and accuracy thresholds, we discard a total of 40 annotations and leave 1,281 activity logs for analysis.
4.2.2 Tool Analysis Results
To study the impact of different tools on our metrics, we use the participants’ activity logs to group these annotations by which tool was used for a majority of pixels. We can then compare these groups’ accuracies and annotation time by using Welch’s t-test, which computes whether the differences in the metrics are significant for groups of unequal variances and sizes. We create four groups based on tool usage as described below:
@@ 4.2.2 Tool Analysis Results
  • Elevation Guided: Annotations where elevation-guided tools (segmentation and both BFS selections) combined were used to label $>50\%$ of all labeled pixels. This group is used to directly compare elevation-guided tools against the brush
Brush BFS Segment. Elevation Guided
@@
Table 1: Mean accuracies and p-values from t-tests against the brush group. We find all three groups that rely on elevation-guided tools to be more accurate than the group using the manual tool. The $\alpha$ needed for significance (computed by BH) for this family of tests was 0.033. All comparisons are statistically significant except the BFS group, likely because the group had a much smaller sample size.
Figure 6: Aggregated annotations from brush and segmentation groups on a study quadrant, visualized using the flattened aggregate view with a certainty threshold of 0.33. In green, notice how the annotations from the brush group do not label the border pixels between the flooded and dry areas, while the annotations from the segmentation group correctly assign labels to these. Also, in pink, notice how the brush group leaves the selected tree canopy area unlabeled, while the segmentation group labels it flooded.
In order, these groups ended up with sizes of 787, 134, 346, and 494 annotations. We find that, even though the instructional video emphasized that participants should rely mostly on the elevation-guided tools, users gravitated towards the more simple and familiar brush tool. This likely shows an example of algorithm aversion [16, 15], where users are proven to prefer human decision-making rather than algorithmic tools even when presented with algorithms that outperform humans. We acknowledge that this could be fixed in future deployments by forcing some usage of the elevation-guided tools, but we leave more in-depth approaches to overcoming algorithm aversion as future work.
@@
We provide the mean accuracies for each group in Table 2. Because we test whether elevation-guided tools increase accuracy, we also present the p-values for their differences against the brush group. From these results, we find that those who relied on elevation-guided tools were able to produce more accurate annotations than those who relied on the brush tool. This shows that elevation-guided tools are effective for improving accuracy (and thus training data quality), even for untrained crowdworkers. We additionally checked for differences between the BFS, segmentation, and elevation-guided groups and found no statistically significant results.
To gain insight into the significance of accuracy differences between groups and explain how elevation-guided tools improve accuracy, we present a qualitative analysis of aggregated annotations from the brush and segmentation groups in Figure 6. For the chosen quadrant, there were 25 annotations in the brush group and 18 in the segmentation group. This example shows how the segmentation tool improves participant annotations by labeling border pixels and tree canopy. This is beneficial for dependent ML models, as elevation border pixels are especially important for elevation-guided training.
Brush BFS Segment. Elevation Guided
Mean 79.61 77.73 48.07 56.74
P-Value vs Brush - 0.434 0.000 0.000
P-Value vs BFS 0.434 - 0.007 0.036
P-Value vs Seg. 0.000 0.007 - 0.129
Table 2: Mean time taken and p-values from t-tests between each group. The $\alpha$ needed for significance (computed by BH) for this family of tests was 0.025. All comparisons are statistically significant except brush against BFS, and elevation-guided against BFS and segmentation. These results show that the segmentation tool greatly decreases annotation time.
@@
Figure 7: Box and whisker plot showing the time taken for annotations in the brush, BFS, segmentation, and elevation-guided groups. We find use of the segmentation tool drastically reduces the time taken to complete an annotation, with brush and BFS groups having median times of 39.97 and 38.18 minutes while segmentation and elevation-guided groups have median times of only 17.45 and 21.65 minutes.
@@
To test annotation speed of the tools, we provide the mean time taken for each group as well as the p-value for their differences against each other in Table 2. From this table, we find that the segmentation tool was significantly faster than the other two tools. To better quantify the tools’ differences, we also present annotation speed results in a box and whisker plot in Figure 7. This plot shows that annotations relying on the segmentation tool were much faster than those that did not, with the median time taken of the segmentation group being 2.29×2.29$\times$2.29 × and 2.19×2.19$\times$2.19 × smaller than the medians of the brush and BFS groups respectively. These results clearly support our segmentation tool as a method of quick annotation while preserving accuracy, and show it outperforms the BFS tool in efficiency. Importantly, these results show that even untrained crowdworkers can use this tool effectively.
@@ 4.3 Crowdsourced Annotations for ML Model Training
In this section, we show how FloodTrace can enable researchers to use crowdsourced annotations as high-quality training data through our aggregation and correction process, which we detail in subsubsection 4.3.1. We propose this step to replace the extremely time-consuming annotation step of researchers’ workflows. To assess the training data quality of the results from aggregation and correction, we use study participants’ annotations as training data for ML models and compare them against models trained with reference annotations subsubsection 4.3.2.
@@ 4.3.1 Example Aggregation and Correction
We conducted an example of our aggregation and correction process on two randomly selected regions using all the annotations for those regions from our user study. We purposefully did not use any accuracy, time, or tool usage cutoffs when creating these sets of annotations in order to ensure there would be inaccuracies in the aggregate data. For each region, there was an average of 45 annotations per pixel. We input these sets to the application frontend, and one of our trained users utilized our aggregate and uncertainty views to correct each region. On average, $4\%$ of each region’s pixels were corrected. This process was timed, and it took an average of 24 minutes to make each region’s correction annotation. We compare this to the average time it took our trained users to create reference annotations using our application, which was 116 minutes per region. As training robust flood detection models requires many different annotated regions, this shows how the crowdsourcing and correction workflow can save researchers huge amounts of time compared to the typical annotation process.
@@ 4.3.2 Prediction Setup and Results
In order to use our crowdsourced annotations (corrected and uncorrected) to train our ML models, we pre-process them by turning each group of annotations into a training set of soft labels. This is done by computing the flood and dry scores for each pixel, where the flood score is the sum of the number of times it was annotated as flooded divided by the number of times it was annotated as either flooded or dry, with dry scores using the same formula. We ignore unlabeled pixels in each annotation to compute these scores. We experimented with binarizing training sets by giving hard labels of flood and dry to pixels with scores above a certainty threshold, but found better performance by training using these soft labels. This is likely because soft labels with scored values are able to encode the probability of pixels being flood or dry, which allows the model to learn more effectively.
@@
Table 3: Performance metrics for models trained using uncorrected, corrected, and reference labels as ground truths. Results show that corrected crowdsourced labels lead to around equal accuracy and f-score as fully expert-created labels, while requiring much less work on the part of the researcher.
We initialized three models, using the architecture described, and trained each of them in the same two regions. Each model used a different set of labels as its respective ground truths during training, with one model using the uncorrected crowdsourced labels, one using the corrected crowdsourced labels, and one using our reference labels. We then use these three trained models to predict flooded and dry labels for an unseen test region and compute our quality metrics. We show these results for each group in Table 3. We find that, with the corrected crowdsourced labels, we are able to achieve around equal results (measured in accuracy and f-score) to our expert-labeled references and improve dramatically over the uncorrected labels.
@@ 5 Conclusion and Limitations
We have presented FloodTrace, a web-based application for quick and accurate annotation of flooded regions that better enables crowdsourcing. This framework was built directly from requirements gathered from researchers in flood extent mapping. Our application utilizes a region’s elevation data to provide users insight in 3D and guide users with semi-automatic annotation tools. This application has already been utilized for the creation of training data by domain experts for cutting-edge flood extent mapping models [54]. Our application brings to the web environment, traditionally complex segmentation methods using topological data analysis to greatly outperform the state-of-the-art elevation-guided tool in efficiency. Our experimental user study shows that the benefits of our tools can be found even for untrained users. With our method for aggregating and correcting groups of annotations, researchers are able to use crowdsourced annotations as training data for their ML models with equal performance to fully expert-labeled data. This improves their workflows greatly, replacing an extremely time-consuming annotation process with a much quicker correction step.
@@
Our work does have limitations. Our work only supports labeling for two classes of pixels, flooded or dry. Selecting features of the data with more specific labels, such as permanent water bodies, is useful for some flood detection models, and future work could tackle extending elevation-guided annotation tools to tasks with more informative labels. In addition, there are limitations with regards to our use of DEM data for our evaluation. The use of DEM data means that objects above the ground level are absent. For some regions, such as heavily urban environments, the presence of these objects could affect water propagation enough to cause the semi-automatic tools to label things incorrectly, for example causing the BFS to falsely label downstream pixels as flooded when the water actually stopped at a building. This limitation could be addressed by using digital surface model (DSM) data or models built from LiDAR data, which take into account above-ground structures. Our application can support any form of elevation data as long as it corresponds to the input satellite imagery, but as of yet we have only evaluated the use of DEM data. Experimentation with other elevation models is left for future work. Finally, as with the machine learning algorithms that rely on elevation data, our system requires that the elevation data and flooded aerial imagery being annotated are collected at similar dates. If not, deviations between elevation at the time of flooding and time of data collection could lead to error.
Acknowledgements.
This work was partly funded by NSF Collaborative Research Awards 2401274, 2221812, and 2106461, and NSF PPoSS Planning and Large awards 2217036 and 2316157.